# Extrapolation for bilinear compact operators in the variable exponent setting

### v1

@@ 2.B. Variable exponent Muckenhoupt characteristics
Given vectors of weights $\vec{w}=(w_{1},\dots,w_{m})$, $\vec{w}_{0}=(w_{0,1},\dots,w_{0,m})$ and $\vec{w}_{1}=(w_{1,1},\dots,w_{1,m})$ we define
@@
The following class of fractional-type variable multiple weights was first introduced in [6].
2.7 Definition ([6], Definition 1.3).
Let $\gamma\in[0,m)$ be a constant. Let $p_{j}(\cdot)\in\mathscr{P}$, $j=1,\dots,m$ and let $q\in\mathscr{P}_{0}$ such that $\gamma=\frac{1}{p(\cdot)}-\frac{1}{q(\cdot)}$, where we define $p\in\mathscr{P}_{0}$ by $\frac{1}{p(\cdot)}=\sum_{j=1}^{m}\frac{1}{p_{j}(\cdot)}$. We say that a vector of weights $\vec{w}=(w_{1},\dots,w_{m})$ satisfies the multilinear $\mathcal{A}^{\gamma}_{\vec{p}(\cdot),q(\cdot)}$ condition (or $\vec{w}\in\mathcal{A}^{\gamma}_{\vec{p}(\cdot),q(\cdot)}$) if
$\vec{w}]_{\mathcal{A}^{\gamma}_{\vec{p}(\cdot),q(\cdot)}}:=\sup_{Q}|Q|^{\gamma-m}\|\nu_{\vec{w}}\chi_{Q}\|_{q(\cdot)}\prod_{j=1}^{m}\|w_{j}^{-1}\chi_{Q}\|_{p_{j}^{\prime}(\cdot)}<\infty,$
where the supremum is taken over all cubes $Q\subset\mathbb{R}^{n}$.
2.8 Remark.
If $\gamma=0$, then $\mathcal{A}_{\vec{p}(\cdot),q(\cdot)}^{\gamma}=\mathcal{A}_{\vec{p}(\cdot)}$. This bi/multilinear version of the variable $\mathcal{A}_{\vec{p}(\cdot)}$ condition was first introduced in [33]. Moreover, notice that in the above definitions we can substitute cubes for balls.
@@ 3. Auxiliary results
3.1 Lemma ([12], Proposition 2.18).
Given a measurable function $p(\cdot):\mathbb{R}^{n}\rightarrow(0,\infty)$ such that $p_{+}<\infty$, there holds that $\||f|^{s}\|_{p(\cdot)}=\|f\|^{s}_{sp(\cdot)},$ for all $s>0.$
The following Hölder’s inequality hold in variable Lebesgue spaces.
3.2 Lemma ([12], Corollary 2.30).
@@
We recall the following corollary of [10, Theorem 1.1] as stated in [22, Lemma 2.8].
3.3 Lemma ([10], Corollary 4.3).
Let $r(\cdot)\in LH\cap\mathcal{P}$ with $r_{+}<\infty$. If $u\in\mathcal{A}_{r(\cdot)}$, then there exists an exponent $s>1$ depending only on $n,$ $r(\cdot)$ and $u]_{\mathcal{A}_{r(\cdot)}}$ such that for all $t\in(1,s)$ and for all cubes cubes $Q\subset\mathbb{R}^{n}$,
$|Q|^{-\frac{1}{{tr_{Q}}}}\|u\chi_{Q}\|_{L^{tr(\cdot)}}\lesssim|Q|^{-\frac{1}{r_{Q}}}\|u\chi_{Q}\|_{L^{r(\cdot)}},$
where the implicit constant depends only on $n$, $r(\cdot)$, $u]_{\mathcal{A}_{r(\cdot)}}$ and $1]_{\mathcal{A}_{\widetilde{v}(\cdot)}}$$\infty$, where
$\widetilde{v}(\cdot):=\frac{r(\cdot)}{\frac{1}{t}-\frac{1}{s}}.$
Now, we emphasize that it is possible to characterize the multilinear $\mathcal{A}_{\vec{p}(\cdot),q(\cdot)}$ weights in terms of the $\mathcal{A}_{p(\cdot)}$ condition. In particular, relying on arguments as in [6, Lemma 1.2] and [33, Proposition 4.7] we obtain the following result.
@@
Let $p_{j}\in\mathscr{P}$, $j=1,\dots,m,$ with $\frac{1}{p(\cdot)}=\sum_{j=1}^{m}\frac{1}{p_{j}(\cdot)}$ and $\gamma=\frac{1}{p(\cdot)}-\frac{1}{q(\cdot)}\in[0,m)$. Then $\vec{w}\in\mathcal{A}^{\gamma}_{\vec{p}(\cdot),q(\cdot)}$ if and only if
(3.5) $\begin{cases}w_{j}^{-\frac{1}{m}}\in\mathcal{A}_{mp_{j}^{\prime}(\cdot)},&j=1,\dots,m,\\ \nu_{\vec{w}}^{\frac{1}{m}}\in\mathcal{A}_{mq(\cdot)}.\end{cases}$
@@
Notice that the proof of “only if” direction is contained in [6, Lemma 1.2]. To prove the “if” direction, we assume that (3.5) holds. Then, by Lemma 3.1 and Hölder’s inequality (Lemma 3.2) with exponents
$\frac{1}{\frac{\gamma+m^{2}}{m}(mq)^{\prime}(\cdot)}+\sum_{j=1}^{m}\frac{1}{\frac{\gamma+m^{2}}{m}(mp_{j}^{\prime})^{\prime}(\cdot)}=1$
@@
$\lesssim|Q|^{-\frac{\gamma}{m}-m}\|(\nu_{\vec{w}}^{-\frac{1}{m}})^{\frac{m}{\gamma+m^{2}}}\chi_{Q}\|_{\frac{\gamma+m^{2}}{m}(mq)^{\prime}(\cdot)}^{\frac{\gamma+m^{2}}{m}}\prod_{j=1}^{m}\|(w_{j}^{\frac{1}{m}})^{\frac{m}{\gamma+m^{2}}}\chi_{Q}\|_{\frac{\gamma+m^{2}}{m}(mp_{j}^{\prime})^{\prime}(\cdot)}^{\frac{\gamma+m^{2}}{m}}$
$=|Q|^{-\frac{\gamma}{m}-m}\|\nu_{\vec{w}}^{-\frac{1}{m}}\chi_{Q}\|_{(mq)^{\prime}(\cdot)}\prod_{j=1}^{m}\|w_{j}^{\frac{1}{m}}\chi_{Q}\|_{(mp_{j}^{\prime})^{\prime}(\cdot)}.$
Hence, this and Definition 2.4 of $\mathcal{A}_{p(\cdot),q(\cdot)}^{0}=\mathcal{A}_{p(\cdot)}$ imply that
$\|\nu_{\vec{w}}\chi_{Q}\|_{q(\cdot)}^{\frac{1}{m}}\prod_{j=1}^{m}\|w_{j}^{-1}\chi_{Q}\|_{p_{j}^{\prime}(\cdot)}^{\frac{1}{m}}$ $=\|\nu_{\vec{w}}^{\frac{1}{m}}\chi_{Q}\|_{mq(\cdot)}\prod_{j=1}^{m}\|w_{j}^{-\frac{1}{m}}\chi_{Q}\|_{mp_{j}^{\prime}(\cdot)}$
$\lesssim|Q|^{-\frac{\gamma}{m}-m}\|\nu_{\vec{w}}^{\frac{1}{m}}\chi_{Q}\|_{mq(\cdot)}\|\nu_{\vec{w}}^{-\frac{1}{m}}\chi_{Q}\|_{(mq)^{\prime}(\cdot)}$
$\qquad\times\prod_{j=1}^{m}\|w_{j}^{\frac{1}{m}}\chi_{Q}\|_{(mp_{j}^{\prime})^{\prime}(\cdot)}\|w_{j}^{-\frac{1}{m}}\chi_{Q}\|_{mp_{j}^{\prime}(\cdot)}$
$\lesssim|Q|^{-\frac{\gamma}{m}-m}|Q||Q|^{m}.$
@@
$\bigg(|Q|^{\gamma-m}\|\nu_{\vec{w}}\chi_{Q}\|_{q(\cdot)}\prod_{j=1}^{m}\|w_{j}^{-1}\chi_{Q}\|_{p_{j}^{\prime}(\cdot)}\bigg)^{\frac{1}{m}}\lesssim 1.$
This and Definition 2.7 of $\mathcal{A}_{\vec{p}(\cdot),q(\cdot)}$ allow us to conclude that $\vec{w}\in\mathcal{A}_{\vec{p}(\cdot),q(\cdot)}$, which completes the proof. ∎
@@
If $m=2$ and $\gamma=0$, then Lemma 3.4 is exactly [33, Proposition 4.7].

### v2

@@ A $m$-admissible quadruple $(\vec{p}(\cdot),q(\cdot),\vec{r},s)$ is said to be proper if
Let $(\vec{p}(\cdot),q(\cdot),\vec{r},s)$ be a $m$-admissible quadruple with
$\frac{1}{p(\cdot)}-\frac{1}{q(\cdot)}=\gamma\in[0,\infty).$
We say that a vector of weights $\vec{w}=(w_{1},\dots,w_{m})$ satisfies the $m$-linear $\mathcal{A}_{({\vec{p}(\cdot),q(\cdot)}),(\vec{r},s)}$ condition (or $\vec{w}\in\mathcal{A}_{({\vec{p}(\cdot),q(\cdot)}),(\vec{r},s)}$) if
(2.7) $\vec{w}]_{\mathcal{A}_{(\vec{p}(\cdot),q(\cdot)),(\vec{r},s)}}:=\sup_{Q}|Q|^{\gamma-\left(\frac{1}{r}-\frac{1}{s}\right)}\|\nu_{\vec{w}}\chi_{Q}\|_{\frac{1}{\frac{1}{q(\cdot)}-\frac{1}{s}}}\prod_{j=1}^{m}\|w_{j}^{-1}\chi_{Q}\|_{\frac{1}{\frac{1}{r_{j}}-\frac{1}{p_{j}(\cdot)}}}<\infty,$
where the supremum is taken over all cubes $Q\subset\mathbb{R}^{n}$ and $\nu_{\vec{w}}:=\prod_{j=1}^{m}w_{j}$.
If $\gamma=0$, we denote
$\mathcal{A}_{\vec{p}(\cdot),(\vec{r},s)}:=\mathcal{A}_{({\vec{p}(\cdot),q(\cdot)}),(\vec{r},s)}.$
Moreover, if $r_{j}=1$ and $s=\infty$, then we denote
$\mathcal{A}_{\vec{p}(\cdot),q(\cdot)}:=\mathcal{A}_{({\vec{p}(\cdot),q(\cdot)}),(\vec{r},s)}.$
2.8 Remark.
  1. (1)
    If $r_{j}=1$, $s=\infty$ and $\gamma\in[0,m)$, then $\mathcal{A}_{({\vec{p}(\cdot),q(\cdot)}),(\vec{r},s)}=\mathcal{A}_{\vec{p}(\cdot),q(\cdot)}$. This recovers the multilinear off-diagonal variable weight class $\mathcal{A}_{\vec{p}(\cdot),q(\cdot)}$ that was first introduced in [undefe, Definition 1.3]. Moreover, in the special case that $\gamma=0$, then $\mathcal{A}_{\vec{p}(\cdot),q(\cdot)}=\mathcal{A}_{\vec{p}(\cdot)}$. The latter bi/multilinear diagonal variable weight class $\mathcal{A}_{\vec{p}(\cdot)}$ was first introduced in [undefai].
  2. (2)
    If $m=1$ and $\gamma\in[0,1)$, then $\mathcal{A}_{({\vec{p}(\cdot),q(\cdot)}),(\vec{r},s)}$ by Lemma 2.4 becomes $\mathcal{A}_{(p(\cdot),q(\cdot)),(\vec{\rho},\vec{s})}$ with $\vec{\rho}=(\rho_{1},\rho_{2})=\left(r,\frac{1}{-\gamma+\frac{1}{r}}\right)$ and $\vec{s}=(s_{1},s_{2})=\left(\frac{1}{\gamma+\frac{1}{s}},s\right)$, where the linear limited range variable weight class $\mathcal{A}_{(p(\cdot),q(\cdot)),(\vec{\rho},\vec{s})}$ was first introduced in [undefag, page 6] but the notation stems from the work [undefv, Definition 1.4].
  3. (3)
    By (2) if $m=1$ and $\gamma=0$, then this definition coincides with the linear diagonal/limited range variable weight class $\mathcal{A}_{p(\cdot),(r,s)}$ first introduced in [undefag, Definition 3.4]. If in addition, we assume that $r_{j}=1$ and $s=\infty$, then $\mathcal{A}_{p(\cdot),(r,s)}=\mathcal{A}_{p(\cdot)}$, where the linear diagonal variable weight class $\mathcal{A}_{p(\cdot)}$ was first introduced in [undefi, (1.3) in page 365].
2.9 Definition.
Given variable exponents
$p_{1}(\cdot),\dots,p_{m}(\cdot),\qquad p_{0,1}(\cdot),\dots,p_{0,m}(\cdot)\qquad\text{and}\qquad p_{1,1}(\cdot),\dots,p_{1,m}(\cdot)$
and vectors of weights
$\vec{w}=(w_{1},\dots,w_{m}),\qquad\vec{w}_{0}=(w_{0,1},\dots,w_{0,m})\qquad\text{and}\qquad\vec{w}_{1}=(w_{1,1},\dots,w_{1,m})$
we define
$\vec{p}(\cdot):=(p_{1}(\cdot),\dots,p_{m}(\cdot)),\quad\vec{p}_{0}(\cdot):=(p_{0,1}(\cdot),\dots,p_{0,m}(\cdot))\quad\text{and}\quad\vec{p}_{1}(\cdot):=(p_{1,1}(\cdot),\dots,p_{1,m}(\cdot))$
@@ 3. Auxiliary results
3.1 Lemma ([undefk], Proposition 2.18).
Let $p(\cdot)\in\mathscr{P}_{0}$. Then, it holds that $\||f|^{s}\|_{p(\cdot)}=\|f\|^{s}_{sp(\cdot)},$ for all constants $s\in(0,\infty)$.
The following Hölder’s inequalities hold in variable Lebesgue spaces.
3.2 Lemma ([undefk], Corollary 2.30).
@@
3.3 Lemma ([undefak], Lemma 6.2).
Let $m\geq 2$ and $p(\cdot),p_{1}(\cdot),\dots,p_{m}(\cdot)\in\mathscr{P}_{0}$ satisfy
$\frac{1}{p(\cdot)}=\sum_{j=1}^{m}\frac{1}{p_{j}(\cdot)},$
where $j=1,\dots,m$. Then, for all $f_{j}\in L^{p_{j}(\cdot)}$ and $f_{1}\cdots f_{m}\in L^{p(\cdot)}$ one has
(3.4) $\|f_{1}\cdots f_{m}\|_{L^{p(\cdot)}}\lesssim\|f_{1}\|_{L^{p_{1}(\cdot)}}\cdots\|f_{m}\|_{L^{p_{m}(\cdot)}},$
where the implicit constant depends only on the $p_{j}(\cdot)$ and $m.$
We point out that Hölder’s inequality for two variable exponents can be found in [undefk, Corollary 2.28] (see also [undefv, Remark 2.6]). Moreover, the results [undefk, Corollary 2.28] and [undefak, Lemma 6.2] were originally stated and proved for variable exponents that belong to the set $\mathcal{P}$. However, by tracking their proofs one can see that they also hold for variable exponents that belong to the set $\mathscr{P}_{0}$.
Now, we emphasize that it is possible to characterize the multilinear $\mathcal{A}_{({\vec{p}(\cdot),q(\cdot)}),(\vec{r},s)}$ weights in terms of the $\mathcal{A}_{(p(\cdot),q(\cdot)),(r,s)}$ and $\mathcal{A}_{p(\cdot),(r,s)}$ weight classes. In particular, relying on similar arguments as in [undefe, Lemma 1.2], [undefaf, Proposition 3.1.6] or [undefai, Proposition 4.7] we obtain the following result.
@@
Let $(\vec{p}(\cdot),q(\cdot),\vec{r},s)$ be a $m$-admissible quadruple. Then $\vec{w}\in\mathcal{A}_{({\vec{p}(\cdot),q(\cdot)}),(\vec{r},s)}$ if and only if
(3.6) $\begin{cases}w_{j}\in\mathcal{A}_{p_{j}(\cdot),(r_{j},\sigma_{j})},&\text{with}\quad\frac{1}{\sigma_{j}}=\frac{1}{r_{j}}-\big(\frac{1}{r}-\frac{1}{s}\big)\quad\text{for all}\quad j=1,\dots,m,\\ \nu_{\vec{w}}\in\mathcal{A}_{(p(\cdot),q(\cdot)),(r,s)}.\end{cases}$
@@
Fix $j_{0}\in\{1,\dots,m\}$. Notice that $(p_{j_{0}}(\cdot),p_{j_{0}}(\cdot),r_{j_{0}},\sigma_{j_{0}})$ is a $1$-admissible quadruple. This follows from the fact that $(p_{j_{0}})_{+}<\sigma_{j_{0}}$. Indeed, this inequality is equivalent to $\frac{1}{\sigma_{j_{0}}}<\frac{1}{(p_{j_{0}})_{+}}$. Then, this is further equivalent to
$\frac{1}{r_{j_{0}}}-\frac{1}{r}+\frac{1}{s}<\frac{1}{(p_{j_{0}})_{+}}.$
By rearrangement, this becomes
$\frac{1}{r_{j_{0}}}-\frac{1}{(p_{j_{0}})_{+}}<\frac{1}{r}-\frac{1}{s}.$
We estimate
$\frac{1}{r}-\frac{1}{p(\cdot)}=\frac{1}{r_{j_{0}}}-\frac{1}{p_{j_{0}}(\cdot)}+\sum_{\begin{subarray}{c}j=1\\ j\neq j_{0}\end{subarray}}^{m}\frac{1}{r_{j}}-\frac{1}{p_{j}(\cdot)}\geq\frac{1}{r_{j_{0}}}-\frac{1}{p_{j_{0}}(\cdot)}.$
Hence, taking essential supremum, we obtain
[…]
as wanted.
Similarly, by Lemma 2.4 we have that $(p(\cdot),q(\cdot),r,s)$ is a $1$-admissible quadruple.
Now, we first prove the “only if” direction, that is, $\vec{w}\in\mathcal{A}_{({\vec{p}(\cdot),q(\cdot)}),(\vec{r},s)}\Rightarrow$ (3.6). Notice that
$\frac{1}{p_{j_{0}}(\cdot)}-\frac{1}{\sigma_{j_{0}}}=\frac{1}{p(\cdot)}-\frac{1}{s}+\sum_{\begin{subarray}{c}j=1\\ j\neq j_{0}\end{subarray}}^{m}\frac{1}{r_{j}}-\frac{1}{p_{j}(\cdot)}\qquad\text{and}\qquad\frac{1}{p(\cdot)}=\frac{1}{q(\cdot)}+\gamma.$
Hence, by Hölder’s inequality (Lemma 3.3) with variable exponents, we have
$\big\|w_{j_{0}}\chi_{Q}\big\|_{\frac{1}{\frac{1}{p_{j_{0}}(\cdot)}-\frac{1}{\sigma_{j_{0}}}}}\big\|w_{j_{0}}^{-1}\chi_{Q}\big\|_{\frac{1}{\frac{1}{r_{j_{0}}}-\frac{1}{p_{j_{0}}(\cdot)}}}$ $=\Bigg\|\nu_{\vec{w}}\prod_{\begin{subarray}{c}j=1\\ j\neq j_{0}\end{subarray}}^{m}w_{j_{0}}^{-1}\chi_{Q}\Bigg\|_{\frac{1}{\frac{1}{p(\cdot)}-\frac{1}{s}+\sum_{\begin{subarray}{c}j=1\\ j\neq j_{0}\end{subarray}}^{m}\frac{1}{r_{j}}-\frac{1}{p_{j}(\cdot)}}}\Big\|w_{j_{0}}^{-1}\chi_{Q}\Big\|_{\frac{1}{\frac{1}{r_{j_{0}}}-\frac{1}{p_{j_{0}}(\cdot)}}}$
$\lesssim\big\|\nu_{\vec{w}}\chi_{Q}\big\|_{\frac{1}{\frac{1}{p(\cdot)}-\frac{1}{s}}}\prod_{j=1}^{m}\big\|w_{j}^{-1}\chi_{Q}\big\|_{\frac{1}{\frac{1}{r_{j}}-\frac{1}{p_{j}(\cdot)}}}$
$\lesssim|Q|^{\gamma}\big\|\nu_{\vec{w}}\chi_{Q}\big\|_{\frac{1}{\frac{1}{q(\cdot)}-\frac{1}{s}}}\prod_{j=1}^{m}\big\|w_{j}^{-1}\chi_{Q}\big\|_{\frac{1}{\frac{1}{r_{j}}-\frac{1}{p_{j}(\cdot)}}}.$
Multiplying both sides of the previous estimate by $|Q|^{-(\frac{1}{r}-\frac{1}{s})}$ and taking a supremum over all cubes $Q$ we get that
$w_{j_{0}}]_{\mathcal{A}_{p_{j_{0}}(\cdot),(r,s)}}\lesssim[\vec{w}]_{\mathcal{A}_{(\vec{p}(\cdot),q(\cdot)),(\vec{r},s)}}.$
For the claim that $\nu_{\vec{w}}\in\mathcal{A}_{(p(\cdot),q(\cdot)),(r,s)}$, first notice that
$\frac{1}{r}-\frac{1}{p(\cdot)}=\sum_{j=1}^{m}\frac{1}{r_{j}}-\frac{1}{p_{j}(\cdot)}.$
Thus, by Hölder’s inequality (Lemma 3.3) with variable exponents, we obtain
$\big\|\nu_{\vec{w}}\chi_{Q}\big\|_{\frac{1}{\frac{1}{q(\cdot)}-\frac{1}{s}}}\big\|\nu_{\vec{w}}^{-1}\chi_{Q}\big\|_{\frac{1}{\frac{1}{r}-\frac{1}{p(\cdot)}}}\lesssim\big\|\nu_{\vec{w}}\chi_{Q}\big\|_{\frac{1}{\frac{1}{q(\cdot)}-\frac{1}{s}}}\prod_{j=1}^{m}\big\|w_{j}^{-1}\chi_{Q}\big\|_{\frac{1}{\frac{1}{r_{j}}-\frac{1}{p_{j}(\cdot)}}}.$
Multiplying both sides of this estimate by $|Q|^{\gamma-(\frac{1}{r}-\frac{1}{s})}$ and taking a supremum over all cubes $Q$ we deduce that
$\nu_{\vec{w}}]_{\mathcal{A}_{(p(\cdot),q(\cdot)),(r,s)}}\lesssim[\vec{w}]_{\mathcal{A}_{(\vec{p}(\cdot),q(\cdot)),(\vec{r},s)}}.$
To prove the “if” direction, we assume that (3.6) holds. Then, by Lemma 3.1 and Hölder’s inequality (Lemma 3.2) with variable exponents
$\frac{\frac{1}{r}-\frac{1}{p(\cdot)}}{m(\frac{1}{r}-\frac{1}{s})}+\sum_{j=1}^{m}\frac{\frac{1}{p_{j}(\cdot)}-\frac{1}{\sigma_{j}}}{m(\frac{1}{r}-\frac{1}{s})}=1$
@@
$\lesssim|Q|^{-m(\frac{1}{r}-\frac{1}{s})}\Bigg\|(\nu_{\vec{w}}^{-1})^{\frac{1}{m(\frac{1}{r}-\frac{1}{s})}}\chi_{Q}\Bigg\|_{\frac{m(\frac{1}{r}-\frac{1}{s})}{\frac{1}{r}-\frac{1}{p(\cdot)}}}^{m(\frac{1}{r}-\frac{1}{s})}\prod_{j=1}^{m}\Bigg\|w_{j}^{\frac{1}{m(\frac{1}{r}-\frac{1}{s})}}\chi_{Q}\Bigg\|_{\frac{m(\frac{1}{r}-\frac{1}{s})}{\frac{1}{p_{j}(\cdot)}-\frac{1}{\sigma_{j}}}}^{m(\frac{1}{r}-\frac{1}{s})}$
$=|Q|^{-m(\frac{1}{r}-\frac{1}{s})}\|\nu_{\vec{w}}^{-1}\chi_{Q}\|_{\frac{1}{\frac{1}{r}-\frac{1}{p(\cdot)}}}\prod_{j=1}^{m}\|w_{j}\chi_{Q}\|_{\frac{1}{\frac{1}{p_{j}(\cdot)}-\frac{1}{\sigma_{j}}}}.$
Hence, this implies that
$\|\nu_{\vec{w}}\chi_{Q}\|_{\frac{1}{\frac{1}{q(\cdot)}-\frac{1}{s}}}\prod_{j=1}^{m}\|w_{j}^{-1}\chi_{Q}\|_{\frac{1}{\frac{1}{r_{j}}-\frac{1}{p_{j}(\cdot)}}}$ $\lesssim|Q|^{-m(\frac{1}{r}-\frac{1}{s})}\|\nu_{\vec{w}}\chi_{Q}\|_{\frac{1}{\frac{1}{q(\cdot)}-\frac{1}{s}}}\|\nu_{\vec{w}}^{-1}\chi_{Q}\|_{\frac{1}{\frac{1}{r}-\frac{1}{p(\cdot)}}}$
$\qquad\times\prod_{j=1}^{m}\|w_{j}\chi_{Q}\|_{\frac{1}{\frac{1}{p_{j}(\cdot)}-\frac{1}{\sigma_{j}}}}\|w_{j}^{-1}\chi_{Q}\|_{\frac{1}{\frac{1}{r_{j}}-\frac{1}{p_{j}(\cdot)}}}$
$\lesssim|Q|^{-m(\frac{1}{r}-\frac{1}{s})}|Q|^{-\gamma+\frac{1}{r}-\frac{1}{s}}|Q|^{m(\frac{1}{r}-\frac{1}{s})}.$
@@
$|Q|^{\gamma-(\frac{1}{r}-\frac{1}{s})}\|\nu_{\vec{w}}\chi_{Q}\|_{\frac{1}{\frac{1}{q(\cdot)}-\frac{1}{s}}}\prod_{j=1}^{m}\|w_{j}^{-1}\chi_{Q}\|_{\frac{1}{\frac{1}{r_{j}}-\frac{1}{p_{j}(\cdot)}}}\lesssim 1.$
Taking a supremum over all cubes $Q$ we conclude that $\vec{w}\in\mathcal{A}_{({\vec{p}(\cdot),q(\cdot)}),(\vec{r},s)}$, which completes the proof. ∎
@@
In view of the following Lemma 3.8, notice that Lemma 3.5 extends [undefe, Lemma 1.2] and [undefai, Proposition 4.7]. In particular, our result deals with variable weight classes that contain the linear/multilinear full range diagonal/off-diagonal weight classes of [undefe, Lemma 1.2].
3.8 Lemma.
Let $(\vec{p}(\cdot),\vec{q}(\cdot),\vec{r},s)$ be a $m$-admissible quadruple. Let also $a\in(0,\infty)$. Then, we have
$\vec{w}\in\mathcal{A}_{(\vec{p}(\cdot),q(\cdot)),(\vec{r},s)}\Leftrightarrow\vec{w}^{\frac{1}{a}}\in\mathcal{A}_{(a\vec{p}(\cdot),aq(\cdot)),(a\vec{r},as)}.$
Moreover, in this case there holds
$\vec{w}^{\frac{1}{a}}]_{\mathcal{A}_{(a\vec{p}(\cdot),aq(\cdot)),(a\vec{r},as)}}=[\vec{w}]_{\mathcal{A}_{(\vec{p}(\cdot),q(\cdot)),(\vec{r},s)}}^{\frac{1}{a}}.$
Proof.
The proof follows immediately by recalling (2.7) and using Lemma 3.1. Indeed, we have
$|Q|^{\frac{\gamma}{a}-\left(\frac{1}{ar}-\frac{1}{as}\right)}\|\nu_{\vec{w}}^{\frac{1}{a}}\chi_{Q}\|_{\frac{1}{\frac{1}{aq(\cdot)}-\frac{1}{as}}}\prod_{j=1}^{m}\|w_{j}^{-\frac{1}{a}}\chi_{Q}\|_{\frac{1}{\frac{1}{ar_{j}}-\frac{1}{ap_{j}(\cdot)}}}$
$\qquad=\left(|Q|^{\gamma-\left(\frac{1}{r}-\frac{1}{s}\right)}\|\nu_{\vec{w}}\chi_{Q}\|_{\frac{1}{\frac{1}{q(\cdot)}-\frac{1}{s}}}\prod_{j=1}^{m}\|w_{j}^{-1}\chi_{Q}\|_{\frac{1}{\frac{1}{r_{j}}-\frac{1}{p_{j}(\cdot)}}}\right)^{\frac{1}{a}}.$
Taking a supremum over all cubes $Q$ we obtain the desired conclusion. ∎